What is , -5/8y = -1/2
13/7n = -13/14
-7/9s= -1/3

Answer by stanbon(75887)

What is ,
-5/8y = -1/2
y = (-1/2)(-6/5)
y = 6/10
y = 3/5
=======================

13/7n = -13/14
n = (-13/14)(7/13)
n = -7/14
n = -1/2
=======================
-7/9s= -1/3
s = (-1/3)(-9/7)
s = 3/7
----------------
Cheers,
Stan H.
=================